# Progressive Compositionality in Text-to-Image Generative Models

### v1

@@ Abstract
Despite the impressive text-to-image (T2I) synthesis capabilities of diffusion models, they often struggle to understand compositional relationships between objects and attributes, especially in complex settings. Existing solutions have tackled these challenges by optimizing the cross-attention mechanism or learning from the caption pairs with minimal semantic changes. However, can we generate high-quality complex contrastive images that diffusion models can directly discriminate based on visual representations? In this work, we leverage large-language models (LLMs) to compose realistic, complex scenarios and harness Visual-Question Answering (VQA) systems alongside diffusion models to automatically curate a contrastive dataset, ConPair, consisting of 15k pairs of high-quality contrastive images. These pairs feature minimal visual discrepancies and cover a wide range of attribute categories, especially complex and natural scenarios. To learn effectively from these error cases, i.e., hard negative images, we propose EvoGen, a new multi-stage curriculum for contrastive learning of diffusion models. Through extensive experiments across a wide range of compositional scenarios, we showcase the effectiveness of our proposed framework on compositional T2I benchmarks. The project page can be found at https://github.com/evansh666/EvoGen.
@@ 1 Introduction
The rapid advancement of text-to-image generative models (Saharia et al., 2022; Ramesh et al., 2022) has revolutionized the field of image synthesis, driving significant progress in various applications such as image editing (Brooks et al., 2023; Zhang et al., 2024), video generation (Brooks et al., 2024) and medical imaging (Han et al., 2024a). Despite their remarkable capabilities, state-of-the-art models such as Stable Diffusion (Rombach et al., 2022) and DALL-E 3 (Betker et al., 2023) still face challenges with composing multiple objects into a coherent scene (Huang et al., 2023; Liang et al., 2023; Majumdar et al., 2024). Common issues include incorrect attribute binding, miscounting, and flawed object relationships as shown in Figure 1. For example, when given the prompt “a red motorcycle and a yellow door”, the model might incorrectly bind the colors to the objects, resulting in a yellow motorcycle.
Recent progress focuses on optimizing the attention mechanism within diffusion models to better capture the semantic information conveyed by input text prompts (Agarwal et al., 2023; Chefer et al., 2023; Pandey et al., 2023). For example, Meral et al. (2023) proposes manipulating the attention on objects and attributes as contrastive samples during test-time to optimize model performance. While more focused, the practical application of these methods still falls short of fully addressing attribute binding and object relationships. Other works advocate to develop compositional generative models to improve compositional performance as each constituent model captures distributions of an independent domain (Du & Kaelbling, 2024). However, such approach assumes a fixed prespecified structure to compose models, limiting generalization to new distributions.
@@
In this paper, we argue that curriculum training is crucial to equip diffusion models with a fundamental understanding of compositionality. Given that existing models often struggle with even basic tasks (e.g., generating three cats when prompted with “Two cats are playing”(Wang et al., 2024), we progressively introduce more complex compositional scenarios during fine-tuning. This staged training strategy helps models build a solid foundation before tackling intricate cases before improving their performance on a wide range of compositional tasks.
Although many datasets exist for compositional generation (Wang et al., 2023; Feng et al., 2023), there remains a significant gap in datasets that offer a clear progression from simple to complex samples within natural and reasonable contexts. Moreover, creating high-quality contrastive image datasets is both costly and labor-intensive, especially given the current limitations of generative models in handling compositional tasks. To address this, we propose an automatic pipeline to generate faithful contrastive image pairs, which we find crucial for guiding models to focus on compositional discrepancies. In summary, our work can be summarized as follows:
@@
We introduce ConPair, a meticulously crafted compositional dataset consisting of high-quality contrastive images with minimal visual representation differences, covering a wide range of attribute categories. By leveraging LLMs, we scale up the complexity of compositional prompts while maintaining a natural context design. Our dataset features faithful images generated by diffusion models, assisted by VQA systems to ensure accurate alignment with the text prompts.
@@
We are the first work to incorporate curriculum contrastive learning into a diffusion model to improve compositional understanding. The process is broken into three streamlined sub-tasks: (1) learning single object-attribute composition, (2) mastering attribute binding between two objects, and (3) handling complex scenes with multiple objects. We conduct extensive experiments using the latest benchmarks and demonstrate that EvoGen significantly boosts the model’s compositional understanding, outperforming most baseline generative methods.
@@ 2.2 Compositional Datasets and Benchmarks
The most commonly used data sets for object-attribute binding, including DrawBench (Saharia et al., 2022), CC-500 (Feng et al., 2023) and Attend-and-Excite (Chefer et al., 2023) construct text prompts by conjunctions of objects and a few of common attributes like color and shape. To more carefully examine how generative models work on each compositional category, recent work explores the disentanglement of different aspects of text-to-image compositionality. Huang et al. (2023) introduces T2I-CompBench that constructing prompts by LLMs which covers six categories including color, shape, textual, (non-)spatial relationships and complex compositions; Recently, Gen-AI (Li et al., 2024a) collects prompts from professional designers which captures more enhanced reasoning aspects such as differentiation, logic and comparison.
Another line of work proposes contrastive textual benchmarks to evaluate the compositional capability of generative models. ABC-6K (Feng et al., 2023) contains contrast pairs by either swapping the order objects or attributes while they focus on negative text prompts with minimal changes. WinogroundT2I (Zhu et al., 2023) contains 11K complex, high-quality contrastive sentence pairs spanning 20 categories. However, such benchmarks focus on text perturbations but do not have images, which have become realistic with the advancement of generative models.
Several benchmarks featuring contrastive image pairs have also been introduced. Compositional Splits C-CUB and C-Flowers (Park et al., 2021) mainly focused on the color and shape attributes of birds and flowers, sourcing from Caltech-UCSD Birds (Wah et al., 2011), Oxford-102 (Flowers) (Nilsback & Zisserman, 2008). Thrush et al. (2022) curated Winoground consists of 400 high-quality contrastive text-image examples. EqBen (Wang et al., 2023) is an early effort to use Stable Diffusion to synthesize images to evaluate the equivariance of VLMs similarity, but it lacks more complex scenarios. Yuksekgonul et al. (2023) emphasizes the importance of hard negative samples and constructs negative text prompts in ARO by swapping different linguistic elements in the captions sourced from COCO and sampling negative images by the nearest-neighbor algorithm. However, it is not guaranteed the negative images found in the datasets truly match the semantic meaning of the prompts.
@@
Three square clock. A brown dog are eating pears to the left of two fluffy cats.
@@ 3.2 Generating contrastive images
Figure 2: EvoGen Framework. Data generation pipeline (left) and curricumlum contrastive learning (right). Quality control of image geneartion (bottom): Given a prompt, SD3 generates multiple candidate images, which are evaluated by LLaVA. We select the best image by alignment and CLIPScore. If the alignment score is low, we prompt LLaVA to describe the image as a new revised caption based on the generated image.
@@
The high-level objective of ConPair is to generate positive images that faithfully adhere to the positive text guidance, while the corresponding negative images do not align with the positive text, despite having minimal visual differences from the positive images. As the quality of images generated by diffusion-based T2I generative models varies significantly (Karthik et al., 2023), we first generate 10-20 candidate images per prompt. However, how to select the most faithful image is difficult. Existing automatic metrics like CLIPScore are not always effective at comparing the faithfulness of images when they are visually similar. To address this, we propose decomposing each text prompt into a set of questions using an LLM and leverage the capabilities of VQA models to rank candidate images by their alignment score, as illustrated in Figure 2 (bottom)  [^1] Examples of decomposed questions are provided in the Appendix A.3. Note the correct answers can be directly extracted from the prompts. Intuitively, we consider an image a success if all the answers are correct or if the alignment is greater than $\theta_{\text{align}}$ for certain categories, such as Complex. After getting aligned images, we select the best image by automatic metric (e.g., CLIPScore).
@@
Given each positive sample, we generate 20 negative images and select the one with the highest similarity to the corresponding positive image, ensuring that the changes between the positive and negative image pairs are minimal. In case of color and texture, we use image editing rather than generation, as it delivers better performance for these attributes. Han et al. (2024b) proposes that human feedback plays a vital role in enhancing model performance. For quality assurance, 3 annotators randomly manually reviewed the pairs in the dataset and filtered 647 pairs that were obviously invalid.
@@ 4 EvoGen: Curriculum Contrastive Fine-tuning
In the first stage, the samples consist of a single object with either a specific attribute (e.g., shape, color, quantity, or texture), a specific action, or within a simple static scene. The differences between the corresponding negative and positive images are designed to be clear and noticeable. For instance, “A man is walking” vs. “A man is eating”, where the actions differ significantly, allowing the model to easily learn to distinguish between them.
@@
For the contrastive objective, we utilize a variant of the InfoNCE loss (van den Oord et al., 2019), which is widely used in contrastive learning frameworks. This loss function is designed to maximize the similarity between the positive image and its corresponding text prompt, while minimizing the similarity between the negative image and the same text prompt. The loss for a positive-negative image pair is expressed as follows:
@@ 5.1 Implementation Details
DALL-E 3 (Betker et al., 2023) 77.85 62.05 70.36 28.65 30.03 37.73
SD v2.1 (Rombach et al., 2022) 50.65 42.21 49.22 13.42 30.96 33.86
EvoGen-SD v2.1 (Ours) 71.04 54.57 72.34 21.76 33.08 42.52
SD3-Medium (Esser et al., 2024) 80.63 58.48 75.24 34.88 31.17 47.72
EvoGen- SD3-Medium (Ours) 83.08 62.83 77.59 36.95 33.12 49.07
Table 3: Alignment evaluation on T2I-CompBench. The best results are in bold
Basic Advanced
Method Attribute Scene Relation Avg Count Differ Compare Logical Avg
Spatial Action Part Negate Universal
SD v2.1 0.75 0.77 0.72 0.72 0.69 0.74 0.66 0.63 0.61 0.50 0.57 0.58
SD-XL Turbo 0.79 0.82 0.77 0.78 0.76 0.79 0.69 0.65 0.64 0.51 0.57 0.60
SD-XL 0.82 0.84 0.80 0.81 0.81 0.82 0.71 0.67 0.64 0.49 0.57 0.60
SD3-Medium 0.86 0.86 0.87 0.86 0.88 0.86 0.74 0.77 0.72 0.50 0.73 0.68
DALL-E 3 0.91 0.91 0.89 0.89 0.91 0.90 0.78 0.76 0.70 0.46 0.65 0.65
Table 4: Gen-AI Benchmark Results.
Figure 4: Average CLIP image-text similarities between the text prompts and the images generated by different models. The Full Prompt Similarity considers full text prompt. Minimum Object represents the minimum of the similarities between the generated image and each of the two object prompts. Example of this benchmark is in Appendix C.
@@
In an attempt to evaluate the faithfulness of generated images, we use GPT-4 to decompose a text prompt into a pair of questions and answers, which serve as the input of our VQA model, LLaVA v1.5 (Liu et al., 2024). Following previous work (Huang et al., 2023; Feng et al., 2023), we finetune and evaluate EvoGen on SD v2.1 (Rombach et al., 2022) and SD3-Medium (Esser et al., 2024).
@@
We compare our results with several state-ofthe-art methods, including trending open-sourced T2I models that trained on large training data, SD v1.4, SD v2.1 and SD3-Medium, DALL-E 2 (Ramesh et al., 2022) and SDXL (Podell et al., 2023). ComposableDiffusion v2 (Liu et al., 2023) is designed for conjunction and negation of concepts for pretrained diffusion models. StructureDiffusion v2 (Feng et al., 2023), Divide-Bind (Li et al., 2024b) and Attn-Exct v2 (Chefer et al., 2023) are designed for attribute binding for pretrained diffusion models. GORs (Huang et al., 2023) finetunes SD v2.1 with selected samples and rewards. PixArt-$\alpha$ (Chen et al., 2023) incorporates cross-attention modules into the Diffusion Transformer. MARS (He et al., 2024) adapts from auto-regressive pre-trained LLMs for T2I generation tasks.
@@
To quantitatively assess the efficacy of our approach, we comprehensively evaluate our method via three primary metrics: 1) compositionality on T2I-CompBench (Huang et al., 2023), 2) GenAI-Bench Li et al. (2024a) and 3) color-object compositionality prompts (Chefer et al., 2023).
@@ 5.2 Performance Comparison and Analysis
Figure 5: Average CLIP similarity of image-text pairs in ConPair.
@@
Beyond the above evaluation, we also assess the alignment between the generated images using EvoGen and text condition on T2I-Compbench. As shown in Table 3, EvoGen achieves SOTA performance across all categories. In particular, we observe significant improvements in the Non-spatial and Complex categories. We hypothesize this is primarily attributed to Stage-III training, where high-quality contrastive samples with complicated compositional components are leveraged to achieve superior alignment capabilities. We further evaluate EvoGen on the Gen-AI benchmark. As indicated in Table 4, EvoGen performs best on all the Advanced prompts, although it exhibits relatively weaker performance in some of the basic categories compared to DALL-E 3.
@@
ConPair + Contra. Loss + Multi-stage FT (EvoGen-SD v2.1) 71.04 54.57 72.34 21.76 33.08 42.52
Table 5: Ablation on T2I-CompBench. ConPair refers to directly finetune SD v2.1 on ConPair.
Figure 6: Qualitative comparison between EvoGen and other SOTA T2I modes with different prompts.
@@
We conduct ablation studies on T2I-CompBench by exploring three key design choices. First, we assess the effectiveness of our constructed dataset, ConPair, by fine-tuning SD v2.1 directly using ConPair. As shown in Table 5, our results consistently outperform the baseline evaluation on SD v2.1 across all categories, demonstrating that our data generation pipeline is effective. Next, we validate the impact of our contrastive loss by comparing it with fine-tuning without this loss. The contrastive loss improves performance in the attribute binding category, though it has less impact on object relationships and complex scenes. We hypothesize this is because attribute discrepancies are easier for the model to detect, while relationship differences are more complex. Finally, applying the multi-stage fine-tuning strategy leads to further improvements, particularly in the Complex category, suggesting that building a foundational understanding of simpler cases better equips the model to handle more intricate scenarios.
@@
Figure 6 presents a side-by-side comparison between EvoGen and other state-of-the-art T2I models, including SDXL, DALL-E 3, SD v3 and PixArt-$\alpha$. EvoGen consistently outperforms the other models in generating accurate images based on the given prompts. SDXL frequently generates incorrect actions and binds attributes to the wrong objects. DALL-E 3 fails to correctly count objects in two examples and misses attributes in the first case. SD v3 struggles with counting and attribute binding but performs well in generating actions. PixArt-$\alpha$ is unable to handle attributes, spatial relationships, and fails to count objects accurately in the second prompt.
@@
Figure 8: User study on 100 randomly selected prompts from Feng et al. (2023). The ratio values indicate the percentages of participants preferring the corresponding model.

### v2

@@ Abstract
Despite the impressive text-to-image (T2I) synthesis capabilities of diffusion models, they often struggle to understand compositional relationships between objects and attributes, especially in complex settings. Existing approaches through building compositional architectures or generating difficult negative captions often assume a fixed prespecified compositional structure, which limits generalization to new distributions. In this paper, we argue that curriculum training is crucial to equipping generative models with a fundamental understanding of compositionality. To achieve this, we leverage large-language models (LLMs) to automatically compose complex scenarios and harness Visual-Question Answering (VQA) checkers to automatically curate a contrastive dataset, ConPair, consisting of 15k pairs of high-quality contrastive images. These pairs feature minimal visual discrepancies and cover a wide range of attribute categories, especially complex and natural scenarios. To learn effectively from these error cases (i.e., hard negative images), we propose EvoGen, a new multi-stage curriculum for contrastive learning of diffusion models. Through extensive experiments across a wide range of compositional scenarios, we showcase the effectiveness of our proposed framework on compositional T2I benchmarks. The project page with data, code, and demos can be found at https://evansh666.github.io/EvoGen_Page/.
@@ 1 Introduction
The rapid advancement of text-to-image generative models (Saharia et al., 2022; Ramesh et al., 2022) has revolutionized the field of image synthesis, driving significant progress in various applications such as image editing (Brooks et al., 2023; Zhang et al., 2024; Gu et al., 2024), video generation (Brooks et al., 2024), and medical imaging (Han et al., 2024a). Despite their remarkable capabilities, state-of-the-art models such as Stable Diffusion (Rombach et al., 2022) and DALL-E 3 (Betker et al., 2023) still face challenges with composing multiple objects into a coherent scene (Huang et al., 2023; Liang et al., 2024b; Majumdar et al., 2024). Common issues include incorrect attribute binding, miscounting, and flawed object relationships as shown in Figure 1. For example, when given the prompt “a red motorcycle and a yellow door”, the model might incorrectly bind the colors to the objects, resulting in a yellow motorcycle.
Recent progress focuses on optimizing the attention mechanism within diffusion models to better capture the semantic information conveyed by input text prompts (Agarwal et al., 2023; Chefer et al., 2023; Pandey et al., 2023). For example, Meral et al. (2023) proposes manipulating the attention on objects and attributes as contrastive samples during the test time to the optimize model performance. While more focused, the practical application of these methods still falls short of fully addressing attribute binding and object relationships. Other works develop compositional generative models to improve compositional performance, as each constituent model captures the distributions of an independent domain (Du & Kaelbling, 2024). However, these approaches assume a fixed prespecified structure to compose models, limiting generalization to new distributions.
@@
In this paper, we argue that curriculum training is crucial to equip diffusion models with a fundamental understanding of compositionality. Given that existing models often struggle with even basic tasks (e.g., generating three cats when prompted with “Two cats are playing”(Wang et al., 2024), we progressively introduce more complex compositional scenarios during fine-tuning. This staged training strategy helps models build a solid foundation before tackling intricate cases for best performance.
With the increasing demand for large-scale data in both model pre-training and fine-tuning, high-quality data generation plays a crucial role in this process (Peng et al., 2025; Ye et al., 2024; Peng et al., 2024a). Although many datasets exist for compositional generation (Wang et al., 2023; Feng et al., 2023a), there remains a significant gap in datasets that offer a clear progression from simple to complex samples within natural and reasonable contexts. Moreover, creating high-quality contrastive image datasets is both costly and labor-intensive, especially given the current limitations of generative models in handling compositional tasks. To address this, we propose an automatic pipeline to generate faithful contrastive image pairs, which we find crucial for guiding models to focus on compositional discrepancies. In summary, our work can be summarized as follows:
@@
We introduce ConPair, a meticulously crafted compositional dataset consisting of high-quality contrastive images with minimal visual representation differences, covering a wide range of attribute categories. By leveraging LLMs, we scale up the complexity of compositional prompts while maintaining a natural context design. Our dataset features faithful images generated by diffusion models, assisted by a VQA checker to ensure accurate alignment with the text prompts.
@@
We also incorporate curriculum contrastive learning into a diffusion model to improve compositional understanding. This curriculum is designed with three sub-tasks: (1) learning single object-attribute composition, (2) mastering attribute binding between two objects, and (3) handling complex scenes with multiple objects. We conduct extensive experiments using the latest benchmarks and demonstrate that EvoGen significantly boosts the model’s compositional understanding, outperforming most baseline generative methods.
@@ 2.2 Compositional Datasets and Benchmarks
The most commonly used data sets for object-attribute binding, including DrawBench (Saharia et al., 2022), CC-500 (Feng et al., 2023a) and Attend-and-Excite (Chefer et al., 2023) construct text prompts by conjunctions of objects and a few of common attributes like color and shape. To more carefully examine how generative models work on each compositional category, recent work explores the disentanglement of different aspects of text-to-image compositionality. Huang et al. (2023) introduces T2I-CompBench that constructing prompts by LLMs which covers six categories including color, shape, textual, (non-)spatial relationships and complex compositions; Recently, Gen-AI (Li et al., 2024a) collects prompts from professional designers which captures more enhanced reasoning aspects such as differentiation, logic and comparison.
Another line of work proposes contrastive textual benchmarks to evaluate the compositional capability of generative models. ABC-6K (Feng et al., 2023a) contains contrast pairs by either swapping the order of objects or attributes while they focus on negative text prompts with minimal changes. WinogroundT2I (Zhu et al., 2023) contains 11K complex, high-quality contrastive sentence pairs spanning 20 categories. However, such benchmarks focus on text perturbations but do not have images, which have become realistic with the advancement of generative models.
Several benchmarks featuring contrastive image pairs have also been introduced. Compositional Splits C-CUB and C-Flowers (Park et al., 2021) mainly focused on the color and shape attributes of birds and flowers, sourcing from Caltech-UCSD Birds (Wah et al., 2011), Oxford-102 (Flowers) (Nilsback & Zisserman, 2008). Thrush et al. (2022) curated Winoground consists of 400 high-quality contrastive text-image examples. EqBen (Wang et al., 2023) is an early effort to use Stable Diffusion to synthesize images to evaluate the equivariance of VLMs similarity, but it lacks more complex scenarios. Yuksekgonul et al. (2023) emphasizes the importance of hard negative samples and constructs negative text prompts in ARO by swapping different linguistic elements in the captions sourced from COCO and sampling negative images by the nearest-neighbor algorithm. However, it is not guaranteed that the negative images found in the datasets truly match the semantic meaning of the prompts.
@@
A brown dog are eating pears to the left of two fluffy cats.
@@ 3.2 Generating contrastive images
Figure 2: EvoGen Framework. Data generation pipeline (left) and curriculum contrastive learning (right). Quality control of image generation (bottom): Given a prompt, SD3 generates multiple candidate images, which are evaluated by LLaVA. We select the best image by alignment and CLIPScore. If the alignment score is low, we prompt LLaVA to describe the image as a newly revised caption based on the generated image.
@@
The high-level objective of ConPair is to generate positive images that faithfully adhere to the positive text guidance, while the corresponding negative images do not align with the positive text, despite having minimal visual differences from the positive images. As the quality of images generated by diffusion-based T2I generative models varies significantly (Karthik et al., 2023), we first generate 10-20 candidate images per prompt. However, selecting the most faithful image is difficult. Existing automatic metrics like CLIPScore are not always effective at comparing the faithfulness of images when they are visually similar. To address this, we propose decomposing each text prompt into a set of questions using an LLM and leverage the capabilities of VQA models to rank candidate images by their alignment score, as illustrated in Figure 2 (bottom)  [^1] Examples of decomposed questions are provided in the Appendix A.3. Note that the correct answers can be directly extracted from the prompts. Intuitively, we consider an image a success if all the answers are correct or if the alignment is greater than $\theta_{\text{align}}$ for certain categories, such as Complex. After getting aligned images, we select the best image by automatic metric (e.g., CLIPScore).
@@
Given each positive sample, we generate 20 negative images and select the one with the highest similarity to the corresponding positive image, ensuring that the changes between the positive and negative image pairs are minimal. In the case of color and texture, we use image editing rather than generation, as it delivers better performance for these attributes. Han et al. (2024b) proposes that human feedback plays a vital role in enhancing model performance. For quality assurance, 3 annotators randomly manually reviewed the pairs in the dataset and filtered 647 pairs that were obviously invalid.
@@ 4 EvoGen: Curriculum Contrastive Fine-tuning
In the first stage, the samples consist of a single object with either a specific attribute (e.g., shape, color, quantity, or texture), a specific action, or a simple static scene. The differences between the corresponding negative and positive images are designed to be clear and noticeable. For instance, “A man is walking” vs. “A man is eating”, where the actions differ significantly, allowing the model to easily learn to distinguish between them.
@@
For the contrastive objective, we utilize a variant of the InfoNCE loss (van den Oord et al., 2019), which is widely used in contrastive learning frameworks. This loss function is designed to maximize the similarity between the positive image and its corresponding text prompt while minimizing the similarity between the negative image and the same text prompt. The loss for a positive-negative image pair is expressed as follows:
@@ 5.1 Implementation Details
EvoGen (Ours) 71.040.13 54.570.25 72.340.26 21.760.18 33.080.35 42.520.38
Table 3: Alignment evaluation on T2I-CompBench. We report average and standard deviations across three runs. The best results are in bold.
@@
In an attempt to evaluate the faithfulness of generated images, we use GPT-4 to decompose a text prompt into a pair of questions and answers, which serve as the input of our VQA model, LLaVA v1.5 (Liu et al., 2024a). Following previous work (Huang et al., 2023; Feng et al., 2023a), we evaluate EvoGen on Stable Diffusion v2 (Rombach et al., 2022).
@@
We compare our results with several state-of-the-art methods, including trending open-sourced T2I models that trained on large training data, Stable Diffusion v1.4 and Stable Diffusion v2 (Rombach et al., 2022), DALL-E 2 (Ramesh et al., 2022) and SDXL (Podell et al., 2023). ComposableDiffusion v2 (Liu et al., 2023) is designed for conjunction and negation of concepts for pretrained diffusion models. StructureDiffusion v2 (Feng et al., 2023a), Divide-Bind (Li et al., 2024b) and Attn-Exct v2 (Chefer et al., 2023) are designed for attribute binding for pretrained diffusion models. GORs (Huang et al., 2023) finetunes Stable Diffusion v2 with selected samples and rewards. PixArt-$\alpha$ (Chen et al., 2023) incorporates cross-attention modules into the Diffusion Transformer. MARS (He et al., 2024) adapts from auto-regressive pre-trained LLMs for T2I generation tasks.
Figure 4: Average CLIP image-text similarities between the text prompts and the images generated by different models. The Full Prompt Similarity considers full-text prompt. Minimum Object represents the minimum of the similarities between the generated image and each of the two object prompts. An example of this benchmark is in subsection C.3.
@@
To quantitatively assess the efficacy of our approach, we comprehensively evaluate our method via two primary metrics: 1) compositionality on T2I-CompBench (Huang et al., 2023)  [^2] More details about specific metrics used in T2I-CompBench are in the Appendix. and 2) color-object compositionality prompts (Chefer et al., 2023).
@@ 5.2 Performance Comparison and Analysis
Figure 5: Average CLIP similarity of image-text pairs in ConPair. Applying VQA checker consistently improves text-image alignment.
@@
Beyond the above evaluation, we also assess the alignment between the generated images using EvoGen and text condition on T2I-Compbench. As depicted in Table 3, we evaluate several crucial aspects, including attribute binding, object relationships, and complex compositions. EvoGen exhibits outstanding performance across 5/6 evaluation metrics. The remarkable improvement in Complex performance is primarily attributed to Stage-III training, where high-quality contrastive samples with complicated compositional components are leveraged to achieve superior alignment capabilities.
@@
ConPair + Contra. Loss + Multi-stage FT 71.04 54.57 72.34 21.76 33.08 42.52
Table 4: Ablation on T2I-CompBench. ConPair refers to directly finetune SDv2 on ConPair.
Figure 6: Qualitative comparison between EvoGen and other SOTA T2I models. EvoGen shows consistent capabilities in following compositional instructions to generate images.
@@
We conduct ablation studies on T2I-CompBench by exploring three key design choices. First, we assess the effectiveness of our constructed dataset, ConPair, by fine-tuning Stable Diffusion v2 directly using ConPair. As shown in Table 4, our results consistently outperform the baseline evaluation on Stable Diffusion v2 across all categories, demonstrating that our data generation pipeline is effective. Next, we validate the impact of our contrastive loss by comparing it with fine-tuning without this loss. The contrastive loss improves performance in the attribute binding category, though it has less impact on object relationships and complex scenes. We hypothesize this is because attribute discrepancies are easier for the model to detect, while relationship differences are more complex. Finally, applying the multi-stage fine-tuning strategy leads to further improvements, particularly in the Complex category, suggesting that building a foundational understanding of simpler cases better equips the model to handle more intricate scenarios.
@@
Figure 6 presents a side-by-side comparison between EvoGen and other state-of-the-art T2I models, including SDXL, DALL-E 3, SD v3 and PixArt-$\alpha$. EvoGen consistently outperforms the other models in generating accurate images based on the given prompts. SDXL frequently generates incorrect actions and binds attributes to the wrong objects. DALL-E 3 fails to correctly count objects in two examples and misses attributes in the first case. SD v3 struggles with counting and attribute binding but performs well in generating actions. PixArt-$\alpha$ is unable to handle attributes and spatial relationships and fails to count objects accurately in the second prompt.
@@
Figure 8: User study on 100 randomly selected prompts from Feng et al. (2023a). The ratio values indicate the percentages of participants preferring the corresponding model.